# A lower bound on the number of bent squares

### v1

@@ 4 A lower bound on $b_{n}$
$m_{n}\sim 2\sqrt{\pi}\left(\frac{2^{\frac{n}{2}-1}}{e}\right)^{2^{\frac{n}{2}}+\frac{1}{2}}=2^{\frac{1}{2}n\cdot 2^{\frac{n}{2}}\left(1+O\left(\frac{1}{n}\right)\right)}.$
@@
$s_{n}\leq\left(2^{\frac{n}{2}-1}-1\right)^{2^{\frac{n}{2}-1}-1}=2^{\frac{1}{4}n\cdot 2^{\frac{n}{2}}\left(1+O\left(\frac{1}{n}\right)\right)}.$
@@
distinct ordered pairs $(A,B)\in M_{n}\times M_{n}$ for which AA and $B$ have the same vertical signature. Those pairs can have at most $s_{n}^{2}$ combined horizontal signatures. By Lemma 4.1 again, it follows that there are at least
@@
The estimate 2 is not intended to be best possible; there are several ways to improve it. However, we are not aware of any refinement of the argument that would do more than altering the implied constant in the term $O\left(\frac{1}{n}\right)$ in 1.

### v2

@@ 4 A lower bound on $b_{n}$
$m_{n}\sim 2\sqrt{\pi}\left(\frac{2^{\frac{n}{2}-1}}{e}\right)^{2^{\frac{n}{2}}+\frac{1}{2}}\implies\operatorname{log}_{2}m_{n}=\frac{1}{2}n\cdot 2^{\frac{n}{2}}\left(1+O\left(\frac{1}{n}\right)\right).$
@@
$s_{n}\leq\left(2^{\frac{n}{2}-1}-1\right)^{2^{\frac{n}{2}-1}-1}\implies\operatorname{log}_{2}s_{n}\leq\frac{1}{4}n\cdot 2^{\frac{n}{2}}\left(1+O\left(\frac{1}{n}\right)\right).$
@@
ordered pairs $(A,B)\in M_{n}\times M_{n}$ for which AA and $B$ have the same vertical signature. Those pairs can have at most $s_{n}^{2}$ distinct combined horizontal signatures. By Lemma 4.1 again, it follows that there are at least
@@
The estimate (2) is not intended to be best possible; there are several ways to improve it. However, we are not aware of any refinement of the argument that would do more than altering the implied constant in the term $O\left(\frac{1}{n}\right)$ in (1).